How fast is a 10 kg box going to be moving from rest after 3 seconds if a 200 N force is applied to the left and a 50 N force is applied to the right?


1
Expert's answer
2021-05-04T12:09:23-0400

The resultant force applied to the box is:


"F = 200N-50N = 150N"

The '-' sign is because the forces directed to the left and to the right are opposite. The resultant force is directed to the left, since "200N>50N".

According to the second Newton's law, the acceleration of the box is:


"a = \dfrac{F}{m}"

where "m = 10kg" is the mass of the box. According to the kinematic equation, the velocity of the box is given as follows:


"v = v_0 +at"

where "v_0 = 0" is the initial velocity of the box (equal to 0 since it starts from rest), and "t = 3s" is the time. Thus, obtain:


"v = \dfrac{F}{m}t\\
v = \dfrac{150N}{10kg}\cdot 3s = 45m/s"

Answer. 45 m/s.